1/ Holiday Laboratories purchased a high-speed industrial centrifuge at a cost of $500,000. Shipping costs totaled $19,000. Foundation work to house the centrifuge cost $8,000. An additional water line had to be run to the equipment at a cost of $2,200. Labor and testing costs totaled $6,800. Materials used up in testing cost $3,300. The capitalized cost is:

Multiple Choice

  • $540,200.

  • $529,200.

  • $539,300.

  • $519,000.

2/ The balance sheets of Davidson Corporation reported net fixed assets of $342,000 at the end of 2018. The fixed-asset turnover ratio for 2018 was 5.0, and sales for the year totaled $1,495,000. Net fixed assets at the end of 2017 were:

Multiple Choice

  • $213,000.

  • $299,000.

  • $256,000.

  • None of these answer choices are correct
Answer #1

1. Capitalized costs

= 500,000+19,000+8,000+2,200+6,800+3,300

= 539,300

Option C

2. Fixed asset turnover = Sales / Average fixed assets

Let year end fixed assets be x

1,495,000 / [(342,000+ x) /2] = 5

2,990,000/(342,000 + x) = 5

598,000 = 342,000 + x

x = 256,000

Option C